Mechanical lab 1

Mechanical lab 1

media%2F36c%2F36c660e8-1b0f-4fd4-80a1-11
Answer #1

I) Refer diagram below:

0 2

Let the initial resistance of all the gages and R3 be R ohms

Let the gage factor br Sg

Let the applied voltage br V volts

From the equation of Wheatstone bridge with strain gages

\Delta E=\frac{r}{(1+r)^2}\left ( \frac{\Delta R_1}{R_1}- \frac{\Delta R_2}{R_2}+ \frac{\Delta R_3}{R_3}- \frac{\Delta R_4}{R_4}\right ) V

r=\frac{R_2}{R_1}=\frac{R_3}{R_4}=\frac{R}{R}=1

Since R2 is aligned perpendicular to strain direction

\epsilon_y=-\nu\epsilon_g

\frac{\Delta R_2}{R_2}=-S_g\nu\epsilon_g+S_g\epsilon_{temp}

Since R3 is not loaded with strain gage

\Delta R_3=0

\frac{\Delta R_1}{R_1}=\frac{\Delta R_4}{R_4}=S_g\epsilon_g+S_g\epsilon_{temp}

Hence

\Delta E=\frac{1}{2^2}\left ( S_g\epsilon_g+S_g\epsilon_{temp}+S_g\nu\epsilon_g- S_g\epsilon_{temp}+ 0-S_g\epsilon _g-S_g\epsilon_{temp}\right ) V=\frac{S_g(\nu\epsilon_g-\epsilon_{temp})}{4} V

Here

\epsilon_{temp}= strain induced by temperature

If a single strain gage is used

\Delta E_1=\frac{1}{4}S_g(\epsilon_g+\epsilon_{temp})V

Hence bridge constant is

\kappa=\frac{\Delta E}{\Delta E_1}=\frac{\nu\epsilon_g-\epsilon_{temp}}{\epsilon_g+\epsilon_{temp}}

Neglecting temperature strain

\kappa=\nu

The \Delta E which is a measure of strain is affected by room temperature as is evident from equation for \Delta E.

II) Refer the diagram below:

The connections of gages have been adjusted for best output

Strain at top surface = \epsilon_b

Strain at bottom surface=-\epsilon_b

The voltage equation is (for the modified circuit)

\Delta E=\frac{r}{(1+r)^2}\left ( \frac{\Delta R_1}{R_1}- \frac{\Delta R_3}{R_3}+ \frac{\Delta R_2}{R_2}- \frac{\Delta R_4}{R_4}\right ) V

\frac{\Delta R_1}{R_1}=-S_g\nu\epsilon_b+S_g\epsilon_{temp}

\frac{\Delta R_2}{R_2}=S_g\epsilon_b+S_g\epsilon_{temp}

\frac{\Delta R_3}{R_3}=S_g\nu\epsilon_b+S_g\epsilon_{temp}

\frac{\Delta R_4}{R_4}=-S_g\epsilon_b+S_g\epsilon_{temp}

Hence

\Delta E=\frac{1}{2^2}\left (( -S_g\nu\epsilon_b+S_g\epsilon_{temp})-(S_g\nu\epsilon_b+S_g\epsilon_{temp})+(S_g\epsilon_b+S_g\epsilon_{temp})-(-S_g\epsilon_b+S_g\epsilon_{temp})\right ) V=\frac{S_g(1-\nu)\epsilon_b}{2} V

Neglecting thermal strain

\Delta E_1=\frac{1}{4}S_g\epsilon_bV

Hence

\kappa=2(1-\nu)

Referring to the formula for \Delta E above, the measured voltaga and hence strain is not affected by room temperature.